Journalist Report – February 21st

Crew 205 Crew Journalist Report 21-FEB-2019 Sol 11

Author: Daniel Robson

Title: “The Happy Little Elves – Hard at Work in their Wintery Workshop”

For those regular readers of the exploits and adventures of Crew 205, we are delighted to tell you that the “Great Jigsaw Puzzle of Sol 9 and 10” was eventually completed last night before bed! Good luck to any future crews that open that Pandora’s box…if you can find it.

The morning began with the crew waking up to a fresh, thick covering of Martian white salt deposit. We are certainly living on a White Mars! While not the Antarctic “Mars”, ours is definitely “Mars” but it’s just white today. Over a leisurely breakfast, the crew discussed tasks for the day, including preparing for the end of our mission and the arrival of the replacement crew – only days away!

The outer modules on our base were swept and mopped and all our equipment so perilously shipped from Earth began to be packed away. Back in the Hab many of the crew began preparing our final Mission Summary and the various other documents we intend to fill with the experiences, data, science, stories, and legends of our time here, out on our own little World.

Nathan, the crew’s GreenHab Officer, has now accepted the fate of his second pair of lost headphones and has now moved onto guarding his third pair with the reverence of a Holy Relic. Natalia, returning to her original role of Commander, and Daniel the HSO, both feared they had fallen to the same curse for a few minutes before finding theirs again.

Hannah has once again outdone herself with cooking today, having made an amazing looking spread of pizza’s and chicken and biscuit crackers for lunch, all garnished with basil from the GreenHab, olive oil, and squeezy cheese. Our multinational crew has the best of all worlds, both in terms of countries and planets!

After lunch, the work began in earnest on our Mission Summary. Like the well-oiled machine our team has proven itself to be, we all jumped at our allocated sections, writing up 11 Sol’s worth of information throughout the afternoon. In fact, we were so successful, we’ve actually had to spend some of tonight cutting it back down to length!

Nathan, Veronica, and Daniel also had a chance this afternoon to take some additional outreach photos for their clubs and groups across the world, another fine example of crew camaraderie and bonding after yesterday’s ceremonial t-shirt swap between the University of Leicester and Florida Institute of Technology.

For dinner (continuing the semi-accidental Italian theme of today) we had a scrumptious pesto pasta, again prepared lovingly by Hannah and her brand new homemade chef’s hat. We’ve heard our replacement crew are French so we do hope they approve of the touch.

Maria, the crew journalist has been working incredibly hard at editing and finalizing the crew videos so that we have some to release before our mission home, and also so we can have a group viewing session together on our last night. Her dedication to the art and outreach of science has given the Author a chance to express his more creative side and write this report (although as you may have noted Santa’s little elves don’t make a direct appearance in this issue).

Ghanim the crew astronomer has been compounded by nights of either blizzards and clouds or a clear night with a massive low hanging Moon causing the rolling landscape around our home to glisten from horizon to horizon and frame the stars in an un-Earthly glow. Unfortunately for him and his telescope, such beauty isn’t very good for the delicate optics, and he has been unable to get the observation of the Orion Nebula he has been after for days now.

Tonight (after the jigsaw has been completed and then buried in one of the under-chair storage containers never to challenge our crew again), we intend to try playing a few games such as dominos or possibly have a rematch of Space Chess. As I’m writing this we’re also trying to work out what each crew member’s catchphrase has been for the mission, and while it might not be useful enough for the Mission Summary we’re having fun with it anyway. Seeing as we’ve also now cracked the code for how to view American DVD’s on a European laptop, we may give Apollo 13 another go.

Hoping the World we are coming home too soon are having a pleasant night, and enjoying the weather, however bleak and beautiful it may be.

Ad Astra!

Greenhab Report – February 21st

Crew 205 Green Hab Report – 21-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (0 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.1°C
High: 22.8°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 10 gallons

Water in Blue Tank: ~ 179 gallons

Crops watered at 0900 hours and 1800 hours:


1) 10.00 AM 19.3° C / 30 %
2) 12.15 PM 21.5° C / 28 %
3) 15:45 PM 20.2° C / 38 %
4) 18:15 PM 19.4° C / 38 %

Changes to crops: None.

Narrative: The A. thaliana experiments will be concluded tomorrow with final results available in the mission summary. The tomatoes and sugar snap peas are visibly looking much better, likely due to the decrease in watering frequency and allowing the soil to aerate. The yellow pigmentation is decreasing and they are more stable in their stem structure. 16.4 g of dead biomass was groomed from several of the plants, including the sugar snap peas.

Fallen material was removed from the surface of the soil to prevent the growth of fungi or other contamination. This material was dried out in the Science Dome oven and this data will contribute of our ISR cellulosic ethanol study. 3.2 g of sage and 8.0 g of basil were harvested for soup and pizza. Similar to yesterday, the door remained closed for most of the day due to the low ambient temperatures and snowy weather.

Harvest: 3.2 g sage

8.0 g basil

Support/supplies needed: None.

Operations Report – February 21st

Crew 205 Operations Report 21 – FEB – 2019

SOL: 11

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1700 (21stFeb), turned off at 1100 (21stFeb)

Solar— SOC % – Turned off Gen (21st Feb) 100%, Turned on Gen (21st Feb) 74%

Diesel Reading – 70%

Propane Reading – ~ 30%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 34%, ~ 189 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 179 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01414305 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 75.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 75.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing to report

Sol Summary – February 21st

Crew 205 Sol Summary Report 21- FEB – 2019

Sol: 11

Summary Title: Covered in White

Author’s name: Natalia Larrea

Mission Status: Nominal

Sol Activity Summary: Today was a quiet day. We woke up surrounded by a white Martian landscape. The snowfall from last night left us with approximately 10 cm of snow on the ground. The Martian weather has been challenging our mission during the past days but despite those challenges, we have been advancing well in our research projects.

It has been 11 days since we arrived to these lands. We can see the end of our mission rising on the horizon. As such, the crew focused today on wrapping up the different projects, start the deep cleaning of the RAM, Science Dome and GreenHab. We also started drafting our mission summary and filmed some last outreach videos. As usual, we have enjoyed a fantastic lunch prepared by our “Martian Chef de Cuisine” Hannah which mastered some home-made Martian Pizza.

Look Ahead Plan: Tomorrow we will perform a deep cleaning of the Hab to make sure everything is ready for the next crew, finish up our mission summary, take some last official pictures and start packing to come back to “Earth”.

Anomalies in work: None

Weather: Consistent snow during most of the day.

Crew Physical Status: Nominal

EVA: None

Reports to be filed: Sol Summary, Journalist Report, GreenHab Report, Operations Report, EVA request (just kidding 😉 we know it is fully covered in white outside)

Support Requested: Paper towel and Febreze spray will be largely appreciated for deep cleaning of the Hab.

Copyright © The Mars Society. All rights reserved. | Main Site